OpenStudy (anonymous):

solve diff equation dy/dx =(sec^2(y))/(1+x^2)

OpenStudy (lgbasallote):

\[\frac{dy}{dx} = \frac{\sec^2 y}{1 + x^2}\] divide both sides by sec^2 y \[\implies \frac{dy}{\sec^2 y dx} = \frac{1}{1+x^2}\] now multiply both sides by dx \[\implies \frac{dy}{\sec^2 y} = \frac{dx}{1+x^2}\] now you can integrate both sides \[\implies \int \frac{dy}{\sec^2 y} = \int \frac{dx}{1+x^2}\] does that help?

OpenStudy (anonymous):

yes. the Integration is the problem i'm having trouble with

OpenStudy (lgbasallote):

okay first turn dy/ sec^2 y into cos^2 y dy you agree those are the same right?

OpenStudy (lgbasallote):

\[\implies \int \cos^2 y dy = \int \frac{dx}{1+x^2}\]

OpenStudy (anonymous):

yes now i see.

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

what about 1/(1+x^2)

OpenStudy (lgbasallote):

trigonometric substitution

OpenStudy (anonymous):

is it tan?

OpenStudy (lgbasallote):

arctan*

OpenStudy (anonymous):

right

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

that's what I thought. Thanks

OpenStudy (lgbasallote):

welcome